# On the Voigt profile and its dual

### v1

@@ 6 Parameter estimation for the Dual Voigt
Thus, parameter estimation can be performed by first estimating the sample mean and sample variance of $T$, denoted with $T_{1}$ and $t_{2}$, and then mapping back the estimates to $\gamma$ and $\sigma$ via
$\widehat{\gamma}=-T_{1}\cdot(T_{2})^{-1}\qquad\textrm{and}\qquad\widehat{\sigma}^{2}=(T_{2})^{-1}$ (10)
In the first stage, it is necessary to find the estimates of a truncated normal distribution (with known truncation values). This is typically done via grid search using an iterative optimization algorithms (17), e.g., the Nelder-Mead algorithm implemented by function optim in the base R package. We present estimates obtained in this way in Table 1 when the true parameter values are $\gamma=\sigma=1$. We highlight that the absolute bias and standard deviation of the estimates are very similar for parameter values randomly chosen in $[0.2,4]$. Overall, the procedure shows a satisfactory performance. Occasionally, we observed failure to reach the minimum and inconsistent values, particularly with small sample sizes $(n=50,100)$.
@@
Table 1: Average point estimates and standard deviation for parameters of the Dual Voigt(1,1)(1,1) calculated over 100100 simulated datasets. Method of moments (MoM) and maximum likelihood (ML) estimates obtained with the Nelder-Mead algorithm.

### v2

@@ 6 Parameter estimation for the Dual Voigt
where the last two parameters are the mean and variance of the underlying normal distribution. Thus, parameter estimation can be performed by first estimating the last two parameters of $T$ with $e_{1}$ and $e_{2}$, and then mapping back the estimates to $\gamma$ and $\sigma$ via
$\widehat{\gamma}=-e_{1}\cdot(e_{2})^{-1}\qquad\textrm{and}\qquad\widehat{\sigma}^{2}=(e_{2})^{-1}$ (10)
In the first stage, it is necessary to find the parameters of a truncated normal distribution (with known truncation values). Maximum Likelihood (ML) and Method of Moments (MoM) are typically found using iterative optimization algorithms (17), e.g., the Nelder-Mead algorithm implemented by function optim in the base R package. We present estimates obtained in this way in Table 1 when the true parameter values are $\gamma=\sigma=1$. We highlight that the absolute bias and standard deviation of the estimates are very similar for parameter values randomly chosen in $[0.2,4]$. Overall, the procedure shows a satisfactory performance. Occasionally, we observed failure to reach the minimum and inconsistent values, particularly with small sample sizes $(n=50,100)$.
@@
Table 1: Average point estimates and standard deviation for parameters of the Dual Voigt(1,1)(1,1) calculated over 100100 simulated datasets. Method of moments (MoM) and maximum likelihood (ML) estimates obtained with the Nelder-Mead algorithm implemented by function optim in base R package.